OpenStudy (anonymous):

what is the answer to 7.5=log[h]

OpenStudy (jack1):

by log... do you mean ln (natural log) or log (base 10)

OpenStudy (anonymous):

log to the base 10

OpenStudy (jack1):

k then 7.5 = log h 7.5 = log h/ log 10 7.5 x log 10 = log h

OpenStudy (jack1):

7.5 x log 10 = log (10^(7.5)) = log h so exponential both sides = h

OpenStudy (anonymous):

thanks :)

OpenStudy (jack1):

cool, what jew get as ur answer?

OpenStudy (jack1):

log h = 17.3 ish...?

OpenStudy (jack1):

Sorry dude, just to clarify: then 7.5 = log h (base 10) 7.5 = log h/ log 10 (both to base e) 7.5 x log 10 = log h (both to base e) 7.5 x log 10 = log (10^(7.5)) = log h (all to base e)

OpenStudy (jack1):

so your answer should start with a 3 if it starts with a 1, your using the wrong base